OpenStudy (meowlover17):

Which statement is true about the product square root of 2(5 + square root of 8)? A,B,C,D ANSWER CHOICES

OpenStudy (meowlover17):

It is rational and equal to 7. It is rational and equal to 9. It is irrational and equal to 4 + square root of 10. It is irrational and equal to 4 + 5square root of 2.

OpenStudy (meowlover17):

\[\sqrt{2}(5+\sqrt{8})\]

OpenStudy (anonymous):

So , we have : =\[5\sqrt{2}+\sqrt{16}=5\sqrt{2}\pm4\]

OpenStudy (meowlover17):

Yes

OpenStudy (meowlover17):

So is the answer D?

OpenStudy (meowlover17):

It is irrational and equal to \[4+5\]

OpenStudy (meowlover17):

\[4+5\sqrt{2}\]

OpenStudy (anonymous):

I'm not sure about C We have + or- 4, but i think it the right one, because none of the others seems to be so :)

OpenStudy (meowlover17):

So the answer is C? o-o

OpenStudy (anonymous):

oww sorry i meant D

OpenStudy (meowlover17):

Oh its fine, yea i thought the same.

OpenStudy (meowlover17):

I need help with 2 more i'll open up another question if thats fine.

OpenStudy (anonymous):

OK , MEDAL pls , i really need them :)

OpenStudy (meowlover17):

Lol almost forgot, i gotcha

OpenStudy (anonymous):

Thanks !